how would you graph this inequality:
y< -1.5x+2.5

Answer by lwsshak3(11628)

how would you graph this inequality:
y< -1.5x+2.5
**
equation for a straight line: y=mx+b, m=slope, b=y-intercept
y= -1.5x+2.5
slope=-1.5=-3/2
y-intercept=2.5
x-intercept:
-1.5x+2.5=0
x=2.5/1.5=5/3=1-2/3
draw a straight line with slope -3/2, y-intercept at 2.5and x-intercept at 1-2/3
then, shade the area below this line